Cold water (Cp = 4180 J/kg · °C) leading to a shower enters a thin-walled double-pipe counter-flow heat exchanger at 15°C at a rate of 0.25 kg/s and is heated to 45°C by hot water (Cp = 4190 J/kg · °C) that enters at 100°C at a rate of 3 kg/s. If the overall heat transfer coefficient is 950 W/m2 · °C. determine the rate of heat transfer and the heat transfer surface area of the heat exchanger using the ε-NTU method.
Answer #1

Giren Data: Ce=4180 Thege To a : Ce = 4180 This Tc; = 150, meke 0.25 kg/see Thi= 100°c Tego = 450 , Man=3 kg/ see U= 950 m mNo to, effectiveness of his can be calculated as E = Tco-Tu = 45-15 = 0.353 1 Thi-To 100-15 Here, Cmin CE i Coronto Cominore